I'm need the point-slope from for this to figure out the other froms of the equations, but im not sure how to get point-slope from from the following infromation... please help.
Line L passes through (1,-6)with slope 3/2. Write the point-slope form, slope-intercept form, and standard form of an equation for line L.

Answer by checkley71(8403)

Y=mX+b m=3/2 THE SLOPE & b=THE Y INTERCEPT
Y=3/2X+b SUBSTITUTE THE X & Y VALUES TO FIND THE Y INTERCEPT
-6=3/2*1+b
-6=3/2+b
b=-6-1.5
b=-7.5
THUS THIS LINE EQUATION IS
Y=3/2X-7.5